# Consciousness-Inspired Spatio-Temporal Abstraction for Better Generalization

### 04 Feb 2024 v3

@@ Abstract
Inspired by human conscious planning, we propose Skipper, a model-based reinforcement learning agent utilizing spatio-temporal abstractions to generalize learned skills in novel situations. It automatically decomposes the given task into smaller, more manageable subtasks, and hence enables sparse decision-making and focused computation on the relevant parts of the environment. This relies on the extraction of an abstracted proxy problem represented as a directed graph, in which vertices and edges are learned end-to-end from hindsight. Our theoretical analyses provide performance guarantees under appropriate assumptions and establish where our approach is expected to be helpful. Generalization-focused experiments validate Skipper’s significant advantage in zero-shot generalization, compared to existing state-of-the-art hierarchical planning methods.
@@ 1 Introduction
By attending to relevant aspects in both space and time, human conscious planning breaks down long-horizon tasks into more manageable abstract steps, each of which can be narrowed down further. Stemming from consciousness in the first sense (C1) (Dehaene et al., 2020), this type of planning focuses attention on mostly the important decision points (Sutton et al., 1999) and relevant environmental factors linking the decision points (Tang et al., 2020), thus operating abstractly both in time and in space. In contrast, existing Reinforcement Learning (RL) agents either operate solely based on intuition (model-free methods) or are limited to reasoning over mostly relatively shortsighted plans (model-based methods) Kahneman (2017). The intrinsic limitations constrain the application of RL in real-world under a glass ceiling formed by challenges of longer-term generalization, below the level of human conscious reasoning.
Building on our previous work on conscious planning (Zhao et al., 2021), we take inspirations to develop a planning agent that automatically decomposes the complex task at hand into smaller subtasks, by constructing abstract “proxy” problems. A proxy problem is represented as a graph where 1) the vertices consist of states proposed by a generative model, corresponding to sparse decision points; and 2) the edges, which define temporally-extended transitions, are constructed by focusing on a small amount of relevant information from the states, using an attention mechanism. Once a proxy problem is constructed and the agent solves it to form a plan, each of the edges defines a new sub-problem, on which the agent will focus next. This divide-and-conquer strategy allows constructing partial solutions that generalize better to new situations, while also giving the agent flexibility to construct abstractions necessary for the problem at hand. Our theoretical analysis establishes guarantees on the quality of the solution to the overall problem.
We also examine empirically whether out-of-training-distribution generalization can be achieved through our method after using only a few training tasks. We show through detailed controlled experiments that the proposed agent, which we name Skipper, performs mostly significantly better in terms of zero-shot generalization, compared to the baselines and to state-of-the-art Hierarchical Planning (HP) methods (Nasiriany et al., 2019; Hafner et al., 2022).
@@ 2 Preliminaries
Reinforcement Learning & Problem Setting. An RL agent interacts with an environment through a sequence of actions to maximize its cumulative reward. The interaction is usually modeled as a Markov decision process (MDP) $\mathcal{M}\equiv\langle\mathcal{S},\mathcal{A},P,R,d,\gamma\rangle$, where $\mathcal{S}$ and $\mathcal{A}$ are the set of states and actions, $P:\mathcal{S}\times\mathcal{A}\to\text{Dist}(\mathcal{S})$ is the state transition function, $R:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}$ is the reward function, $d:\mathcal{S}\to\text{Dist}(\mathcal{S})$ is the initial state distribution, and $\gamma\in[0,1]$ is the discount factor. The agent needs to learn a policy $\pi:\mathcal{S}\to\text{Dist}(\mathcal{A})$ that maximizes the value function, i.e. the expected discounted cumulative reward $\mathbb{E}_{\pi,P}[\sum_{t=0}^{T_{\perp}}\gamma^{t}R(S_{t},A_{t},S_{t+1})|S_{0}\sim d]$, where $T_{\perp}$ denotes the time step at which the episode terminates. A value estimator $Q:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ can be used to guide the search for a good policy. However, real-world problems are typically partially observable, meaning that at each time step $t$, after taking an action $a_{t}\in\mathcal{A}$, the agent receives an observation $x_{t+1}\in\mathcal{X}$, where $\mathcal{X}$ is the observation space. The agent then needs to infer the state from its sequence of observations, which is usually done through a state encoder.
@@
One important goal of RL is to achieve high (generalization) performance on evaluation tasks after learning from a limited number of training tasks, where the evaluation and training distributions may differ; for instance, a policy for a robot may need to be trained in a simulated environment for safety reasons, but would need to be deployed on a physical device, a setting called sim2real. Discrepancy between task distributions is often recognized as a major reason why RL agents are yet to be applied pervasively in the real world (Igl et al., 2019). To address this issue, in this paper, agents are trained on a small set of fixed training tasks, then evaluated in unseen tasks, where there are environmental variations, but the core strategies needed to finish the task remain consistent, for example because of the existence of causal mechanisms (Zhang et al., 2020). To generalize well, the agents need to build learned skills which capture the consistent knowledge across tasks.
@@
Deep Model-based RL. Deep model-based RL uses predictive or generative models to guide the search for a good policy (Silver et al., 2017). In terms of generalization, rich models, expressed by Neural Networks (NNs), may capture generalizable information and infer latent causal structure. Background planning agents e.g., Dreamer (Hafner et al., 2023) use a model as a data generator to improve the value estimators and policies, which they execute in background while interacting with the environment (Sutton, 1991). These agents do not improve on the trained policy at decision time. In contrast, decision-time planning agents e.g., MuZero (Schrittwieser et al., 2020) and PlaNet (Hafner et al., 2019) actively use models at decision time to make better decisions. Recently, Alver & Precup (2022) suggests that the latter approach provides better generalization, aligning with observations from cognitive behaviors (Momennejad et al., 2017).
@@
Options & Goal-Conditioned RL. Temporal abstraction allows RL agents to use sub-policies, and to model the environment over extended time scales, to achieve both better generalization and solving larger problems. Options and their models provide a formalism for temporal abstraction in RL (Sutton et al., 1999). Each option consists of an initiation condition, a policy, and a termination condition. For any set of options defined on an MDP, the decision process that selects only among those options, executing each to termination, is a Semi-MDP (SMDP) (Sutton et al., 1999; Puterman, 2014), consisting of the set of states $\mathcal{S}$, the set of options $\mathcal{O}$, and for each state-option pair, an expected return, and a joint distribution of the next state and transit time. In this paper, we focus on goal-conditioned options, where the initiation set covers the whole state space $\mathcal{S}$. Each such option is a tuple $o=\langle\pi,\beta\rangle$, where $\pi:\mathcal{S}\to\text{Dist}(\mathcal{A})$ is the (intra-)option policy and $\beta:\mathcal{S}\to\{0,1\}$ indicates when a goal state is reached. Hindsight Experience Replay (HER) (Andrychowicz et al., 2017) is often used to train goal-conditioned options by sampling a transition $\langle x_{t},a_{t},r_{t+1},x_{t+1}\rangle$ together with an additional observation $x^{\odot}$ from the same trajectory, which is re-labelled as a “goal”.
@@ 3 Skipper: Spatially & Temporally Abstract Planning
In this section, we describe the main ingredients of Skipper - an agent that formulates a proxy problem for a given task, solves this problem, and then proceeds to “fill in” the details of the plan.
@@ 3.1 Proxy Problems
Proxy problems are finite graphs constructed at decision-time, whose vertices are states and whose directed edges estimate transitions between the vertices, as shown in Fig. 1. We call the states selected to be vertices of the proxy problems checkpoints, to differentiate from other uninvolved states. The current state is always included as one of the vertices. The checkpoints are proposed by a generative model and represent a subset of states that the agent might experience in the current episode, often denoted as $S^{\odot}$ in the rest of the paper. Each edge is annotated with estimates of the duration and reward associated with the transition between the connected checkpoints; these estimates are learned over the relevant aspects of the environment and depend on the agent’s capability. As the low-level policy implementing checkpoint transitions improves, the edges strengthen. Planning in a proxy problem is temporally abstract, since the checkpoints act as sparse decision points. Estimating each checkpoint transition is spatially abstract, as an option corresponding to such a task would base its decisions only on some aspects of the environment state (Bengio, 2017; Konidaris & Barto, 2009), in order to improve generalization as well as computational efficiency (Zhao et al., 2021).
Figure 1: A Proxy Problem on a Grid-World Navigation Task: the MDP of the original problem is in gray and the terminal states are marked with squares. An agent needs to get from the (filled red) position, to the goal (filled green). The agent can reach a distant goal by leveraging a proxy problem with $12$ checkpoints (outlined orange).
@@
The theorem indicates that once the agent achieves high accuracy estimation of the model for the proxy problem and a near-optimal lower-level policy ${\color[rgb]{1,0,0}\pi}$, it converges toward optimal performance (proof in Appendix J.2). The theorem also makes no assumption on $\pi$ because it would likely be difficult to learn a good $\pi$ for far away targets. Despite the generality, in the experiments, we limit ourselves to navigation tasks with sparse rewards for reaching goals, where the goals are included as permanent vertices in the proxy problems. This is a case where the accuracy assumption can be met non-trivially, i.e., while avoiding degenerate proxy problems whose edges involve no rewards. Following Thm. 1, we train estimators for $v_{\pi}$ and $\gamma_{\pi}$ and refer to this as edge estimation.
@@ 3.2 Design Choices
To plan over proxy problems, our framework embraces the following design choices:
@@
Spatio-temporal abstraction: temporal abstraction breaks down the given task into smaller ones, while spatial abstraction [^1] We use “spatial abstraction” to denote specifically the behavior of constraining decision-making to only the relevant environmental factors during an option. Please check Section 4 for discussions and more details. over the state features improves local learning and generalization;
@@
Delusion suppression: we propose a delusion suppression technique to minimize the behavior of chasing non-existent outcomes. This is done by exposing the edge estimators to targets that would otherwise not exist in experience.
@@ 3.3 Problem 1: Edge Estimation
First, we discuss how to estimate the edges of the proxy problem, given a set of already generated checkpoints. Inspired by conscious information processing in brains (Dehaene et al., 2020) and existing approach in Sylvain et al. (2019), we introduce a local perceptive field selector, $\sigma$, consisting of an attention bottleneck that (soft-)selects the top-$k$ local segments of the full state (e.g. a feature map by a typical convolutional encoder); all segments of the state compete for the $k$ attention slots, i.e. irrelevant aspects of state are discarded, to form a partial state representation (Mott et al., 2019; Tang et al., 2020; Zhao et al., 2021; Alver & Precup, 2023). We provide an example in Fig. 2 (see purple parts). Through $\sigma$, the auxiliary estimators, to be discussed soon, force the bottleneck mechanism to promote aspects relevant to the local estimation of connections between the checkpoints. The rewards and discounts are then estimated from the partial state $\sigma(S)$, based on the agent’s behavior.
@@ 3.3.2 Estimate Connections
We learn the connection estimates with auxiliary reward signals that are designed to be task-invariant (Zhao et al., 2019). These estimates are learned using distributional RL, where the output of each estimator takes the form of a histogram over scalar support (Dabney et al., 2018).
@@
Cumulative Distances / Discounts. The cumulative discount leading to the target $s_{\odot}$ under ${\color[rgb]{1,0,0}\pi}$ is unfortunately more difficult to learn than ${\color[rgb]{0,0.5,0}V_{\pi}}$, since the prediction would be heavily skewed towards $1$ if $\gamma\approx 1$. Yet, we can instead effectively estimate cumulative (truncated) distances (or trajectory length) under $\pi$. Such distances can be learned with policy evaluation, where the auxiliary reward is $+1$ on every transition, except at the targets:
@@
Please refer to the Appendix J.1 for the properties of the learning rules for $\hat{v}_{\pi}$ and $\hat{\gamma}_{\pi}$.
Figure 2: Skipper Framework: 1) Partial states consist of a few local fields, soft-selected via top-$k$ attention (Gupta et al., 2021). Skipper’s edge estimations and low-level behaviors $\pi$ are based on the partial states. 2) The checkpoint generator learns by splitting the full state into context and partial descriptions, and fusing them to reconstruct the input. It generates checkpoints by sampling partial descriptions and combines them with the episodic contexts; 3) We prune the vertices and edges of the denser graphs to extract sparser proxy problems. Once a plan is formed, the immediate checkpoint target is used to condition the policy. In the proxy problem example, blue edges are estimated to be bidirectional and red edges have the other direction pruned.
@@ 3.4 Problem 2: Vertex Generation
The checkpoint generator aims to directly model the possible future states without needing to know how exactly the agent might reach them. The details of checkpoint transitions will be abstracted by the connection estimates instead.
@@
We sample in hindsight a diverse distribution of target encoded (full) states $S^{\odot}$, given any current $S_{t}$. Hence, we use as generator a conditional Variational AutoEncoder (VAE) (Sohn et al., 2015) which learns a distribution $p(S^{\odot}|C(S_{t}))=\sum_{z}{p(S^{\odot}|C(S_{t}),z)p(z|C(S_{t}))}$, where $C(S_{t})$ is the extracted context from $S_{t}$ and $z$s are the partial descriptions. We train the generator by minimizing the evidence lower bound on $\langle S_{t},S^{\odot}\rangle$ pairs chosen with HER.
Similarly to Hafner et al. (2023), we constrain the partial description as a bundle of binary variables and train them with the straight-through gradients (Bengio et al., 2013). These binary latents can be easily sampled or composed for generation. Compared to models such as that in Director (Hafner et al., 2022), which generates intermediate goals given the on-policy trajectory, ours is expected to generate a more diverse distribution of states, beneficial for planning in novel scenarios.

### 16 Mar 2024 v4

# Consciousness-Inspired Spatio-Temporal Abstractions for Better Generalization in Reinforcement Learning

@@ Abstract
Inspired by human conscious planning, we propose Skipper, a model-based reinforcement learning framework utilizing spatio-temporal abstractions to generalize better in novel situations. It automatically decomposes the given task into smaller, more manageable subtasks, and thus enables sparse decision-making and focused computation on the relevant parts of the environment. The decomposition relies on the extraction of an abstracted proxy problem represented as a directed graph, in which vertices and edges are learned end-to-end from hindsight. Our theoretical analyses provide performance guarantees under appropriate assumptions and establish where our approach is expected to be helpful. Generalization-focused experiments validate Skipper’s significant advantage in zero-shot generalization, compared to some existing state-of-the-art hierarchical planning methods.
@@ 1 Introduction
Attending to relevant aspects in both time and space, human conscious planning breaks down long-horizon tasks into more manageable steps, each of which can be narrowed down further. Stemming from consciousness in the first sense (C1) (Dehaene et al., 2020), this type of planning focuses attention on mostly the important decision points (Sutton et al., 1999) and relevant environmental factors linking the decision points (Tang et al., 2020), thus operating abstractly both in time and in space. In contrast, existing Reinforcement Learning (RL) agents either operate solely based on intuition (model-free methods) or are limited to reasoning over mostly relatively shortsighted plans (model-based methods) Kahneman (2017). The intrinsic limitations constrain the real-world application of RL under a glass ceiling formed by challenges of generalization.
Building on our previous work on conscious planning (Zhao et al., 2021), we take inspirations to develop a planning agent that automatically decomposes the complex task at hand into smaller subtasks, by constructing abstract “proxy” problems. A proxy problem is represented as a graph where 1) the vertices consist of states imagined by a generative model, corresponding to sparse decision points; and 2) the edges, which define temporally-extended transitions, are constructed by focusing on a small amount of relevant information from the states, using an attention mechanism. Once a proxy problem is constructed and the agent solves it to form a plan, each of the edges defines a new sub-problem, on which the agent will focus next. This divide-and-conquer strategy allows constructing partial solutions that generalize better to new situations, while also giving the agent flexibility to construct abstractions necessary for the problem at hand. Our theoretical analyses establish guarantees on the quality of the solution to the overall problem.
We also examine empirically advantages of out-of-training-distribution generalization of our method after using only a few training tasks. We show through detailed controlled experiments that the proposed framework, named Skipper, in most cases performs significantly better in terms of zero-shot generalization, compared to the baselines and to some state-of-the-art Hierarchical Planning (HP) methods (Nasiriany et al., 2019; Hafner et al., 2022).
@@ 2 Preliminaries
Reinforcement Learning & Problem Setting. An RL agent interacts with an environment via a sequence of actions to maximize its cumulative reward. The interaction is usually modeled as a Markov Decision Process (MDP) $\mathcal{M}\equiv\langle\mathcal{S},\mathcal{A},P,R,d,\gamma\rangle$, where $\mathcal{S}$ and $\mathcal{A}$ are the set of states and actions, $P:\mathcal{S}\times\mathcal{A}\to\text{Dist}(\mathcal{S})$ is the state transition function, $R:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}$ is the reward function, $d:\mathcal{S}\to\text{Dist}(\mathcal{S})$ is the initial state distribution, and $\gamma\in[0,1]$ is the discount factor. The agent needs to learn a policy $\pi:\mathcal{S}\to\text{Dist}(\mathcal{A})$ that maximizes the value of the states, i.e. the expected discounted cumulative reward $\mathbb{E}_{\pi,P}[\sum_{t=0}^{T_{\perp}}\gamma^{t}R(S_{t},A_{t},S_{t+1})|S_{0}\sim d]$, where $T_{\perp}$ denotes the time step at which the episode terminates. A value estimator $Q:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ can be used to guide the search for a good policy. However, real-world problems are typically partially observable, meaning that at each time step $t$, instead of states, the agent receives an observation $x_{t+1}\in\mathcal{X}$, where $\mathcal{X}$ is the observation space. The agent then needs to infer the state from the sequence of observations, usually with a state encoder.
@@
One important goal of RL is to achieve high (generalization) performance on evaluation tasks after learning from a limited number of training tasks, where the evaluation and training distributions may differ; for instance, a policy for a robot may need to be trained in a simulated environment for safety reasons, but would need to be deployed on a physical device, a setting called sim2real. Discrepancy between task distributions is often recognized as a major reason why RL agents are yet to be applied pervasively in the real world (Igl et al., 2019). To address this issue, in this paper, agents are trained on a small set of fixed training tasks, then evaluated in unseen tasks, where there are environmental variations, but the core strategies needed to finish the task remain consistent. To generalize well, the agents need to build learned skills which capture the consistent knowledge across tasks.
@@
Deep Model-based RL. Deep model-based RL uses predictive or generative models to guide the search for policies (Silver et al., 2017). In terms of generalization, rich models, expressed by Neural Networks (NNs), may capture generalizable information and infer latent causal structure. Background planning agents e.g., Dreamer (Hafner et al., 2023) use a model as a data generator to improve the value estimators and policies, which executes in background without directly engaging in the environment (Sutton, 1991). These agents do not improve on the trained policy at decision time. In contrast, decision-time planning agents e.g., MuZero (Schrittwieser et al., 2020) and PlaNet (Hafner et al., 2019) actively use models at decision time to make better decisions. Recently, Alver & Precup (2022) suggests that the latter approach provides better generalization, aligning with observations from cognitive behaviors (Momennejad et al., 2017).
@@
Options & Goal-Conditioned RL. Temporal abstraction allows agents to use sub-policies, and to model the environment over extended time scales, to achieve both better generalization and the divide and conquer of larger problems. Options and their models provide a formalism for temporal abstraction in RL (Sutton et al., 1999). Each option consists of an initiation condition, a policy, and a termination condition. For any set of options defined on an MDP, the decision process that selects only among those options, executing each to termination, is a Semi-MDP (SMDP) (Sutton et al., 1999; Puterman, 2014), consisting of the set of states $\mathcal{S}$, the set of options $\mathcal{O}$, and for each state-option pair, an expected return, and a joint distribution of the next state and transit time. In this paper, we focus on goal-conditioned options, where the initiation set covers the whole state space $\mathcal{S}$. Each such option is a tuple $o=\langle\pi,\beta\rangle$, where $\pi:\mathcal{S}\to\text{Dist}(\mathcal{A})$ is the (intra-)option policy and $\beta:\mathcal{S}\to\{0,1\}$ indicates when a goal state is reached. Hindsight Experience Replay (HER) (Andrychowicz et al., 2017) is often used to train goal-conditioned options by sampling a transition $\langle x_{t},a_{t},r_{t+1},x_{t+1}\rangle$ together with an additional observation $x^{\odot}$ from the same trajectory, which is re-labelled as a “goal”.
@@ 3 Skipper: Spatially & Temporally Abstract Planning
In this section, we describe the main ingredients of Skipper - a framework that formulates a proxy problem for a given task, solves this problem, and then proceeds to “fill in” the details of the plan.
@@ 3.1 Proxy Problems
Proxy problems are finite graphs constructed at decision-time, whose vertices are states and whose directed edges estimate transitions between the vertices, as shown in Fig. 1. We call the states selected to be vertices of the proxy problems checkpoints, to differentiate from other uninvolved states. The current state is always included as one of the vertices. The checkpoints are proposed by a generative model and represent some states that the agent might experience in the current episode, often denoted as $S^{\odot}$ in this paper. Each edge is annotated with estimates of the cumulative discount and reward associated with the transition between the connected checkpoints; these estimates are learned over the relevant aspects of the environment and depend on the agent’s capability. As the low-level policy implementing checkpoint transitions improves, the edges strengthen. Planning in a proxy problem is temporally abstract, since the checkpoints act as sparse decision points. Estimating each checkpoint transition is spatially abstract, as an option corresponding to such a task would base its decisions only on some aspects of the environment state (Bengio, 2017; Konidaris & Barto, 2009), to improve generalization as well as computational efficiency (Zhao et al., 2021).
Figure 1: A Proxy Problem on a Grid-World Navigation Task: the MDP of the original problem is in gray and the terminal states are marked with squares. An agent needs to get from the (filled red) position, to the goal (filled green). Distant goals can be reached by leveraging a proxy problem with $12$ checkpoints (outlined orange).
@@
The theorem indicates that once the agent achieves high accuracy estimation of the model for the proxy problem and a near-optimal lower-level policy ${\color[rgb]{1,0,0}\pi}$, it converges toward optimal performance (proof in Appendix D.2). The theorem also makes no assumption on $\pi$, since it would likely be difficult to learn a good $\pi$ for far away targets. Despite the theorem’s generality, in the experiments, we limit ourselves to navigation tasks with sparse rewards for reaching goals, where the goals are included as permanent vertices in the proxy problems. This is a case where the accuracy assumption can be met non-trivially, i.e., while avoiding degenerate proxy problems whose edges involve no rewards. Following Thm. 1, we train estimators for $v_{\pi}$ and $\gamma_{\pi}$ and refer to this as edge estimation.
@@ 3.2 Design Choices
To implement planning over proxy problems, Skipper embraces the following design choices:
@@
Spatio-temporal abstraction: temporal abstraction breaks down the given task into smaller ones, while spatial abstraction [^1] We use “spatial abstraction” to denote specifically the behavior of constraining decision-making to the relevant environmental factors during an option. Please check Section 4 for discussions and more details. over the state features improves local learning and generalization;
@@
Delusion suppression: we propose a delusion suppression technique to minimize the behavior of chasing non-existent outcomes. This is done by exposing edge estimation to imagined targets that would otherwise not exist in experience.
@@ 3.3 Problem 1: Edge Estimation
First, we discuss how to estimate the edges of the proxy problem, given a set of already generated checkpoints. Inspired by conscious information processing in brains (Dehaene et al., 2020) and existing approach in Sylvain et al. (2019), we introduce a local perceptive field selector, $\sigma$, consisting of an attention bottleneck that (soft-)selects the top-$k$ local segments of the full state (e.g. a feature map by a typical convolutional encoder); all segments of the state compete for the $k$ attention slots, i.e. irrelevant aspects of state are discouraged or discarded, to form a partial state representation (Mott et al., 2019; Tang et al., 2020; Zhao et al., 2021; Alver & Precup, 2023). We provide an example in Fig. 2 (see purple parts). Through $\sigma$, the auxiliary estimators, to be discussed soon, force the bottleneck mechanism to promote aspects relevant to the local estimation of connections between the checkpoints. The rewards and discounts are then estimated from the partial state $\sigma(S)$, conditioned on the agent’s policy.
@@ 3.3.2 Estimate Connections
We learn the connection estimates with auxiliary reward signals that are designed to be not task-specific (Zhao et al., 2019). These estimates are learned using C51-style distributional RL, where the output of each estimator takes the form of a histogram over scalar support (Dabney et al., 2018).
@@
Cumulative Distances / Discounts. With C51 and uniform supports, the cumulative discount leading to $s_{\odot}$ under ${\color[rgb]{1,0,0}\pi}$ is unfortunately more difficult to learn than ${\color[rgb]{0,0.5,0}V_{\pi}}$, since the prediction would be heavily skewed towards $1$ if $\gamma\approx 1$. Yet, we can instead effectively estimate cumulative (truncated) distances (or trajectory length) under $\pi$. Such distances can be learned with policy evaluation, where the auxiliary reward is $+1$ on every transition, except at the targets:
@@
Please refer to the Appendix D.1 for the properties of the learning rules for $\hat{v}_{\pi}$ and $\hat{\gamma}_{\pi}$.
Figure 2: Skipper Framework: 1) Partial states consist of a few local fields, soft-selected via top-$k$ attention (Gupta et al., 2021). Skipper’s edge estimations and low-level behaviors $\pi$ are based on the partial states. 2) The checkpoint generator learns by splitting the full state into context and partial descriptions, and fusing them to reconstruct the input. It imagines checkpoints by sampling partial descriptions and combining them with the episodic contexts; 3) We prune the vertices and edges of the denser graphs to extract sparser proxy problems. Once a plan is formed, the immediate checkpoint target is used to condition the policy. In the proxy problem example, blue edges are estimated to be bidirectional and red edges have the other direction pruned.
@@ 3.4 Problem 2: Vertex Generation
The checkpoint generator aims to directly imagine the possible future states without needing to know how exactly the agent might reach them nor worrying about if they are reachable. The feasibility of checkpoint transitions will be abstracted by the connection estimates instead.
@@
We sample in hindsight a diverse distribution of target encoded (full) states $S^{\odot}$, given any current $S_{t}$. Hence, we make the generator a conditional Variational AutoEncoder (VAE) (Sohn et al., 2015) which learns a distribution $p(S^{\odot}|C(S_{t}))=\sum_{z}{p(S^{\odot}|C(S_{t}),z)p(z|C(S_{t}))}$, where $C(S_{t})$ is the extracted context from $S_{t}$ and $z$s are the partial descriptions. We train the generator by minimizing the evidence lower bound on $\langle S_{t},S^{\odot}\rangle$ pairs chosen with HER.
Similarly to Hafner et al. (2023), we constrain the partial description as a bundle of binary variables and train them with the straight-through gradients (Bengio et al., 2013). These binary latents can be easily sampled or composed for generation. Compared to models such as that in Director (Hafner et al., 2022), which generates intermediate goals given the on-policy trajectory, ours can generate and handle a more diverse distribution of states, beneficial for planning in novel scenarios.